1. Inputs of a MPa Inputs of b MPb 1 25 1 40 2 20 2...

1.

Inputs of a MPa Inputs of b MPb
1 25 1 40
2 20 2 36
3 15 3 32
4 10 4 24
5 5 5 20
6 2 6 16
7 1 7 8

The table gives marginal product data for resources a and b. The output of these independent resources sells in a purely competitive market at $1 per unit. Assuming the prices of resources a and b are $5 and $8 respectively, what is the least costly combination of resources for the firm to employ in producing 192 units of output?

Group of answer choices

3 of a and 4 of b

2 of a and 6 of b

4 of a and 3 of b

6 of a and 2 of b

2.

Units of Resource 0 1 2 3 4 5
Total Product 0 6 11 15 18 20
Total Revenue ($) 0 36 55 60 54 40

Refer to the table. The marginal revenue product of the third unit of the resource is

Group of answer choices

$3.

$36.

$19.

$5.
Answer #1

Answer

1)

The correct answer is (a) 3 of a and 4 of b.

Cost minimizing condition :

(i) MPa/Pa = MPb/Pb

(ii) Quantity(Q) = 192

Here Pa = 5 and Pb = 8

When Firm hires 3 units of a and 4 units of b then MPa/Pa = 15/5 = 3 and MPb/Pb = 24/8 = 3 and thus condition (i) is satisfied.

When Firm hires 3 units of a and 4 units of b then Total Quantity produced(Q) = 25 + 20 + 15 + 40 + 36 +32 + 24 = 192 units. Thus condition (ii) is also satisfied.

Hence, the correct answer is (a) 3 of a and 4 of b.

2)

The correct answer is (d) $5.

Marginal Revenue Product is the additional revenue received when firm hires 1 additional unit of input(resource).

So, Marginal Revenue Product of 3rd unit = Total Revenue from 3 resources - Total Revenue from 2 resources
= 60 - 55 = $5

Hence, the correct answer is (d) $5.